PPS web services are still experiencing intermittent unavailability and/or slow response times and our staff is working to resolve this problem as soon as possible.
This problem impacts access to various PPS web services:
ftp://arthurhou.pps.eosdis.nasa.gov (GPM FTP data)
http://pps.gsfc.nasa.gov (PPS home page)
https://storm-pps.gsfc.nasa.gov (PPS STORM)
http://registration.pps.eosdis.nasa.gov (PPS GPM registration)
GPM near real time...
Presently, PPS web services are experiencing intermittent unavailability and/or slow response times. PPS is aware of this issue and our System Programmers are working diligently to correct these issues.
The current anomaly affects our arthurhou server which include the PPS Public site http://pps.gsfc.nasa.gov/ , FTP services, STORM: https://storm.pps.eosdis.nasa.gov/ and THOR (THOR online) , etc.
We hope to have all these services restored to normal as soon as possible and will...
Effective about 18 UTC July 21 the GPM MOC stopped sending data to the PPS. This was not a satellite or instrument issue. The problem appeared to be a network connection issue at the Emergency MOC which is currently receiving data as part of a monthly test of the backup facility.
Effective 21:41 UTC the data flow from the MOC to PPS resumed. Apparently there was a major network issue on that part of the network. It has now been fixed and MOC is sending...
PPS is releasing version 3.70.4 of the PPS TKIO toolkit for TRMM/GPM products. This package, sample files, and the supporting documentation can be found at: ftp://gpmweb2.pps.eosdis.nasa.gov/pub/PPStoolkit/GPM/tkio-3.70.4/
Learn more about the TKIO toolkit here: http://pps.gsfc.nasa.gov/ppstoolkit.html
If you have any questions please contact the PPS Helpdesk: firstname.lastname@example.org
Starting at 16:33 UTC July 6 a filled log directory led to sh commands failing on redirecting logs to be created in that directory. A typo in the cleanup routine failed to clean our all the files after the normal 4 days. As a result they have been accumulating since the beginning of the mission. It is unexpected that the sh failed when it could not redirect anymore.
All data products that failed have been created and put in the appropriate directories...
The Tropical Rainfall Measuring Mission (TRMM) spacecraft re-entered the Earth’s atmosphere on June 15, 2015, at 11:55 p.m. EDT, over the South Indian Ocean, according to the U.S. Strategic Command’s Joint Functional Component Command for Space through the Joint Space Operations Center (JSpOC). The U.S. Space Surveillance Network, operated by the Defense Department's JSpOC, had been closely monitoring TRMM’s descent since the mission was ended in April. Most of the spacecraft was expected to...
Starting at 08:38 UTC PPS stopped getting data from the GPM Mission Operations
Center. Data was resumed at 17:21 UTC. However, new GPS data was sent before older GPS data. The science data was sent out of order with the GPS data. This meant that about 125 mins of 1B and 1C GMI data had no geolocation and perhaps more after this had questionable geolocation. The same issues obviously also affected the radar and combined NRT which are just missing for the...
NOAA has reprocessed the global IR data for 10 UTC 26 April
through 14 UTC 27 April due to dropped images, and all 3B41RT and
3B42RT files for this time period have been reprocessed by PPS and are
3B41RT.2015042610.7.bin.gz through 3B41RT.2015042714.7.bin.gz
3B42RT.2015042612.7.bin.gz through 3B42RT.2015042715.7.bin.gz
Such partial dropouts in the IR data result in somewhat lower quality for IMERG Early and Late Runs, but are not cause for reprocessing...
On Sunday, April 26, 2015, between the hours of 12 noon (17 UTC) and 12 midnight EDT (05 UTC April 27) , changes will be implemented to NASA’s Internet access. The purpose of this outage is to complete the Corporate Route Symmetry Project which is being implemented by the Communications Services Office (CSO). As a result of these changes, users can expect an outage to the below services during this time period.
The issues with NOAA's 4-km Merged IR data are closed and the IMERG Early and Late Runs have been restarted from the point at which they stopped. This will provide a continuous record for each, but it also means that it will take a while to process the backlog of data and catch up to the nominal latency.